A tiny sphere of mass 7.40 mg and charge-2.80 nC is initially at a distance of 1.54 μm from a fixed charge of +9.17 nC. (a) If the 7.40-mg sphere is released from rest, find its kinetic energy when it is 0.500 μm from the fixed charge. 0.462 Your response differs from the correct answer by more than 10%. Double check your calculations. J (b) If the 7.40-mg sphere is released from rest, find its speed when it is 0.500 μm from the fixed charge. m/s
Answer #1

a)

using energy conservation

KE = change in PE

KE = k q1 q2 * ( 1/ r1 - 1/r2)

KE = 9*10^9*9.17*10^-18* 2.8* ( 1/ 0.5 - 1/1.54)*10^6

KE = 0.312 J

=======

b)

0.5 m v^2 = 0.312

0.5* 7.4/10^-6 * v^2 = 0.312

v = 290.44 m/s

=======

do comment in case any doubt, will reply for sure.. Goodluck